# Gaussian Process-Based Prediction and Control of Hammerstein-Wiener Systems

### v1

@@ III-C Hyperparameter Tuning
However, the hyperparameters in this problem are high-dimensional, especially since $\left[\Gamma_{1}\ \Gamma_{2}\right]\in\mathbb{R}^{n_{y}L^{\prime}\times(n_{u}L+n_{y}L_{0})}$, which is known to induce overfitting [29]. This problem can be alleviated by selecting a suitable prior distribution for the hyperparameter, known as the hyperprior [30]. Note that $\Gamma_{1}$, $\Gamma_{2}$ correspond to the multi-step-ahead ARX parameters of $G(q)$. The problem of designing priors for linear model parameters has been widely studied in system identification literature (see [3] and references therein). If the state-space parameters are known, we have from [31, Equation 14]
$\Gamma_{1}=:\left[\Gamma_{11}\ \Gamma_{12}\right],\ \Gamma_{11}=\mathcal{O}_{L^{\prime}}\mathcal{P},\ \Gamma_{12}=\mathcal{T}_{L^{\prime}},\ \Gamma_{2}=\mathcal{O}_{L^{\prime}}\mathcal{Q},$
where
$\mathcal{C}_{i}$ $:=\begin{bmatrix}A^{i-1}B&\cdots&AB&B\end{bmatrix}.$
Note that $\Gamma_{12}$ has a block lower-diagonal Toeplitz structure where the last block row corresponds to the system’s Markov parameters of length $L^{\prime}$. For a stable system $G(q)$, a standard stable kernel design such as the TC (tune/correlated) kernel can be applied as the hyperprior of the last block row. There is no explicit structure for $\Gamma_{11}$ and $\Gamma_{2}$. In this work, we follow the same idea in [4, Section 5.3] and model each row of $\Gamma_{11}$ and $\Gamma_{2}$ with a stable kernel indexed by the input-output lag.
@@
$\mathrm{cov}\left(\left(\Gamma_{12}\right)_{i,j},\left(\Gamma_{12}\right)_{i^{\prime},j^{\prime}}\right)$ =$=\begin{cases}0,&j>i\text{ or }j^{\prime}>i^{\prime},\\ s_{i-j,i^{\prime}-j^{\prime}},&\text{otherwise}.\end{cases}$
$\mathrm{cov}\left(\left(\Gamma_{11}\right)_{i,j},\left(\Gamma_{11}\right)_{i^{\prime},j^{\prime}}\right)$ =$=\begin{cases}0,&i\neq i^{\prime},\\ s_{L_{0}+i-j,L_{0}+i^{\prime}-j^{\prime}},&\text{otherwise},\end{cases}$
$\mathrm{cov}\left(\left(\Gamma_{2}\right)_{i,j},\left(\Gamma_{2}\right)_{i^{\prime},j^{\prime}}\right)$ =$=\begin{cases}0,&i\neq i^{\prime},\\ s_{L_{0}+i-j,L_{0}+i^{\prime}-j^{\prime}},&\text{otherwise},\end{cases}$
where $s_{i,j}(\zeta)$ is a stable kernel parameterized by $\zeta$, which is sometimes known as the hyper-hyperparameters. Throughout this work, the TC kernel is used:
$s^{\text{TC}}_{i,j}(\zeta):=\lambda\alpha^{\max{(i,j)}},\ \zeta:=[\lambda\ \alpha]^{\top},\ \lambda\geq 0,\ 0\leq\alpha<1.$
The hyper-hyperparameters $\zeta$ can be selected by cross-validation. For alternative stable kernel design, readers are referred to [3, Section 5.5].
With hyperpriors introduced for $\Gamma_{1}$ and $\Gamma_{2}$, instead of the maximum marginal likelihood problem (13) which considers all hyperparameters as deterministic variables, a joint maximum-a-posteriori/maximum likelihood problem (JMAP-ML) is considered [32]:
$\Theta&=\text{arg}\underset{\Theta}{\text{max}}\ p\left(\boldsymbol{\chi}^{d}|\boldsymbol{\eta}^{d},\boldsymbol{\theta},\Gamma\right)p(\Gamma|\zeta)\\ &=\text{arg}\underset{\Theta}{\text{min}}\ \mathrm{logdet}\left(\Lambda(\Theta)\right)+\mathbf{m}\!^{\top}\!(\boldsymbol{\eta}^{d})\Lambda^{-1}(\Theta)\mathbf{m}(\boldsymbol{\eta}^{d})+\log p(\Gamma|\zeta),\end{aligned}$
@@
$S_{\boldsymbol{\gamma}}:=\mathrm{blkdiag}(S_{L_{0}:1},S_{L_{0}+1:2},\dots,S_{L-1:L^{\prime}},S_{L_{0}:1},S_{L_{0}+1:2},\dots,S_{L-1:L^{\prime}},S_{L^{\prime}-1:0}),$ (28)
@@
Θ¯=argminΘ¯logdet(Λ($\bar{\Theta}=\text{arg}\underset{\bar{\Theta}}{\text{min}}\ \mathrm{logdet}\left(\Lambda(\bar{\Theta})\right)+\mathbf{m}\!^{\top}\!(\boldsymbol{\eta}^{d})\Lambda^{-1}(\bar{\Theta})\mathbf{m}(\boldsymbol{\eta}^{d})+\boldsymbol{\gamma}^{\top}S_{\boldsymbol{\gamma}}^{-1}\boldsymbol{\gamma},$
where $\bar{\Theta}:=\left\{\boldsymbol{\theta},\boldsymbol{\gamma}\right\}$.
Algorithm 1 summarizes the procedure of implicit Gaussian process data-driven prediction for Hammerstein-Wiener systems.
@@
1:Given: hyper-hyperparameters $\zeta$, Hankel data matrices $H_{u}$, $H_{y}$
@@
3:Construct $\Gamma_{1}$, $\Gamma_{2}$ from $\gamma$.
4:Input: Prediction condition $\mathbf{u}$, $\mathbf{y}_{\text{p}}$
5:Solve (24), (25), or (26) for $\mathbf{y}_{\text{f}}$, where $m_{\text{p}}(\eta)$, $k_{\text{p}}(\eta)$ are given by (10), (11), (17)-(22).
6:Output: Output prediction $\mathbf{y}_{\text{f}}$
III-D Recovering Nonlinearities
In addition to output prediction, it could also be of interest to recover the input and output nonlinearities $\psi(\cdot)$ and $\phi(\cdot)$. Note that as discussed in Remark 1, the nonlinearities can only be determined up to scaling factors.
To obtain $\Psi(\mathbf{u})$ at any vector of query points $\mathbf{u}$, the joint distribution of $\Psi(\mathbf{u})$ and $\boldsymbol{\chi}^{d}$ is given by
$\begin{bmatrix}\Psi(\mathbf{u})\\ \boldsymbol{\chi}^{d}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mathbf{m}_{u}(\mathbf{u})\\ \mathbf{m}(\boldsymbol{\eta}^{d})\end{bmatrix},\begin{bmatrix}\mathbf{k}_{u}(\mathbf{u},\mathbf{u})&\boldsymbol{\kappa}_{u}^{\top}(\boldsymbol{\eta}^{d},\mathbf{u})\\ \boldsymbol{\kappa}_{u}(\boldsymbol{\eta}^{d},\mathbf{u})&K(\boldsymbol{\eta}^{d},\boldsymbol{\eta}^{d})+\bar{\Sigma}\end{bmatrix}\right),$
where, similar to the derivation in Section III-B,
$\boldsymbol{\kappa}_{u}(\boldsymbol{\eta}^{d},\mathbf{u}):=\left(\mathbb{I}\otimes\Gamma_{1}\right)\mathbf{k}_{u}(\mathrm{vec}\left(H_{u}\right),\mathbf{u}).$ (30)
Similarly, define
$\boldsymbol{\kappa}_{y}(\boldsymbol{\eta}^{d},\mathbf{y}):=\left(\mathbb{I}\otimes\left[\Gamma_{2}\ -\mathbb{I}\right]\right)\mathbf{k}_{y}(\mathrm{vec}\left(H_{y}\right),\mathbf{y})$ (31)
for $\Phi(\mathbf{y})$. This leads to the posterior distributions of $\Psi(\mathbf{u})$ and $\Phi(\mathbf{y})$:
$\begin{split}\Psi(\mathbf{u})|\boldsymbol{\chi}^{d}&\sim\mathcal{N}\left(\mathbf{m}_{u\text{p}}(\mathbf{u}),\mathbf{k}_{u\text{p}}(\mathbf{u})\right),\\ \Phi(\mathbf{y})|\boldsymbol{\chi}^{d}&\sim\mathcal{N}\left(\mathbf{m}_{y\text{p}}(\mathbf{y}),\mathbf{k}_{y\text{p}}(\mathbf{y})\right),\end{split}$ (32)
@@ III-D Recovering Nonlinearities
$\mathbf{m}_{u\text{p}}(\mathbf{u})$ $:=\mathbf{m}_{u}(\mathbf{u})-\boldsymbol{\kappa}_{u}^{\top}(\boldsymbol{\eta}^{d},\mathbf{u})\left(K+\bar{\Sigma}\right)^{-1}\mathbf{m}(\boldsymbol{\eta}^{d}),$
$\mathbf{k}_{u\text{p}}(\mathbf{u})$ $:=\mathbf{k}_{u}(\mathbf{u},\mathbf{u})-\boldsymbol{\kappa}_{u}^{\top}(\boldsymbol{\eta}^{d},\mathbf{u})\left(K+\bar{\Sigma}\right)^{-1}\boldsymbol{\kappa}_{u}(\boldsymbol{\eta}^{d},\mathbf{u}).$
and similarly for $\mathbf{m}_{y\text{p}}(\mathbf{y})$ and $\mathbf{k}_{y\text{p}}(\mathbf{y})$. Then, $\Psi(\mathbf{u})$ and $\Phi(\mathbf{y})$ are estimated as $\mathbf{m}_{u\text{p}}(\mathbf{u})$ and $\mathbf{m}_{y\text{p}}(\mathbf{y})$, respectively.
IV Data-Driven Predictive Control with Implicit Gaussian Process Model
This section applies receding horizon data-driven predictive control using the implicit Gaussian process predictor described in Section III. In detail, considering the problem of tracking an output reference $r_{k}$ at time $k$, the following expected control cost is considered at time $k$:
$J_{\text{ctr}}:=\left\lVert\mathbf{u}^{k}\right\rVert_{R}^{2}+\mathbb{E}\left[\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{r}^{k})\right\rVert_{Q}^{2}\right],$ (33)
where $R\in\mathbb{R}^{n_{u}L^{\prime}\times n_{u}L^{\prime}}$, $Q\in\mathbb{R}^{n_{y}L^{\prime}\times n_{y}L^{\prime}}$ are the input and the output cost matrices, respectively, and $\mathbf{r}^{k}:=\mathrm{col}\left(r_{k},\dots,r_{k+L^{\prime}-1}\right)$, $\mathbf{u}^{k}:=\mathrm{col}\left(u_{k},\dots,u_{k+L^{\prime}-1}\right)$, $\mathbf{y}^{k}_{0}:=\mathrm{col}\left(y_{k,0},\dots,y_{k+L^{\prime}-1,0}\right)$ are the reference, input, and true output trajectories at time $k$, respectively. In addition, we also consider input constraints $\mathbf{u}^{k}\in\mathcal{U}_{k}$ and linear output constraints $H\mathbf{y}^{k}_{0}\leq q$, where $H\in\mathbb{R}^{n_{c}\times n_{y}L^{\prime}}$ and $q\in\mathbb{R}^{n_{c}}$. Since the prediction error is unbounded, the output constraint can only be enforced with high probability $p$ as a chance constraint $\mathrm{Pr}\left(H\mathbf{y}^{k}_{0}\leq q\right)\geq p$.
In this section, the following assumption is considered.
Assumption 2
@@ IV Data-Driven Predictive Control with Implicit Gaussian Process Model
As discussed in Section III, the distribution of $\Phi(\mathbf{y}^{k}_{0})$ is given by
$\Phi(\mathbf{y}^{k}_{0})\sim\mathcal{N}\left(\Phi(\mathbf{y}^{k})+m_{\text{p}}(\eta^{k}),\hat{k}_{\text{p}}(\eta^{k})\right),$ (34)
where $\mathbf{y}^{k}$ is the output prediction and
@@
Then, we have the following lemma.
Lemma 2
Under Assumptions 1 and 2, the chance constraint $\mathrm{Pr}\left(H\mathbf{y}^{k}_{0}\leq q\right)\geq p$ is guaranteed by

### v2

@@ 3.5 Hyperparameter Tuning
However, the hyperparameters in this problem are high-dimensional, especially since $\left[\Gamma_{1}\ \Gamma_{2}\right]\in\mathbb{R}^{n_{y}L^{\prime}\times(n_{u}L+n_{y}L_{0})}$, which is known to induce overfitting [35]. This problem can be alleviated by selecting a suitable prior distribution for the hyperparameter, known as the hyperprior [18]. Note that $\Gamma_{1}$, $\Gamma_{2}$ correspond to the coefficients of the linear multi-step-ahead predictor for $G(q)$. The problem of designing priors for linear model parameters has been widely studied in the system identification literature. Most prominently, a class of stable spline kernels has been developed; see [31] and references therein.
From (6), it can be seen that $\Gamma_{12}$ has a block lower-diagonal Toeplitz structure where the last block row corresponds to the system’s Markov parameters of length $L^{\prime}$. For a stable system $G(q)$, a standard stable spline kernel design, such as the TC (tune/correlated) kernel, can be applied as the hyperprior of the last block row. There is no explicit structure for $\Gamma_{11}$ and $\Gamma_{2}$. In this work, we follow the same idea as in [32, Section 5.3] and model each row of $\Gamma_{11}$ and $\Gamma_{2}$ with a stable spline kernel indexed by the input-output lag.
@@
$\begin{aligned} \mathrm{cov}\left(\left(\Gamma_{12}\right)_{i,j},\left(\Gamma_{12}\right)_{i^{\prime},j^{\prime}}\right)&=\begin{cases}0,&j>i\ \mathrm{or}\ j^{\prime}>i^{\prime},\\ s_{i-j,i^{\prime}-j^{\prime}},&\mathrm{otherwise}.\end{cases}\\ \mathrm{cov}\left(\left(\Gamma_{11}\right)_{i,j},\left(\Gamma_{11}\right)_{i^{\prime},j^{\prime}}\right)&=\begin{cases}0,&i\neq i^{\prime},\\ s_{L_{0}+i-j,L_{0}+i^{\prime}-j^{\prime}},&\mathrm{otherwise},\end{cases}\\ \mathrm{cov}\left(\left(\Gamma_{2}\right)_{i,j},\left(\Gamma_{2}\right)_{i^{\prime},j^{\prime}}\right)&=\begin{cases}0,&i\neq i^{\prime},\\ s_{L_{0}+i-j,L_{0}+i^{\prime}-j^{\prime}},&\mathrm{otherwise},\end{cases}\end{aligned}$
where $s_{i,j}(\zeta)$ is a stable spline kernel parameterized by $\zeta$, which is sometimes known as the hyper-hyperparameters. Throughout this work, the TC kernel is used:
$s^{\mathrm{TC}}_{i,j}(\zeta):=\lambda\alpha^{\max{(i,j)}},\ \zeta:=[\lambda\ \alpha]^{\top},\ \lambda\geq 0,\ 0\leq\alpha<1.$
The hyper-hyperparameters $\zeta$ can be selected by cross-validation. For alternative stable spline kernel designs, readers are referred to [31, Section 5.5].
With hyperpriors introduced for $\Gamma_{1}$ and $\Gamma_{2}$, instead of the maximum marginal likelihood problem (37) which considers all hyperparameters as deterministic variables, a joint maximum-a-posteriori/maximum-likelihood (JMAP-ML) problem is considered [44]:
$\Theta=\mathrm{arg}\underset{\Theta}{\mathrm{max}}\ q\left(\mathcal{M},\boldsymbol{\chi}^{d}|\boldsymbol{\theta},\Gamma\right)p(\Gamma|\zeta),$ (44)
@@
$S_{\boldsymbol{\gamma}}:=\mathrm{blkdiag}(S_{L_{0}:1},S_{L_{0}+1:2},\dots,S_{L-1:L^{\prime}},\\ S_{L_{0}:1},S_{L_{0}+1:2},\dots,S_{L-1:L^{\prime}},S_{L^{\prime}-1:0}),$ (45)
@@
$\bar{\Theta}=\mathrm{arg}\underset{\bar{\Theta}}{\mathrm{min}}\ \mathrm{logdet}\left(\Lambda_{\mathrm{j}}(\bar{\Theta})\right)+\delta_{\mathrm{j}}^{\top}\Lambda^{-1}_{\mathrm{j}}(\bar{\Theta})\delta_{\mathrm{j}}+\boldsymbol{\gamma}^{\top}S_{\boldsymbol{\gamma}}^{-1}\boldsymbol{\gamma},$ (46)
where $\bar{\Theta}:=\left\{\boldsymbol{\theta},\boldsymbol{\gamma}\right\}$ and $\delta_{\mathrm{j}}$ is defined below (35).
Remark 10.
Solving (46) requires implementing Algorithm 1 inside the optimization problem, which can be time-consuming. For computational efficiency, one can alternatively solve
$\bar{\Theta}=\mathrm{arg}\underset{\bar{\Theta}}{\mathrm{min}}\ \mathrm{logdet}\left(\Lambda(\bar{\Theta})\right)+\delta^{\top}\Lambda^{-1}(\bar{\Theta})\delta+\boldsymbol{\gamma}^{\top}S_{\boldsymbol{\gamma}}^{-1}\boldsymbol{\gamma},$ (47)
where $\delta$ is defined below (15), without the virtual derivative points, and only include them in the posterior prediction.
Algorithm 2 summarizes the procedure of implicit GP data-driven prediction for H-W systems, where lines 2 and 3 estimate the hyperparameters offline and line 5 obtains output predictions online.
@@
1:Given: Hankel data matrices $H_{u}$, $H_{y}$, virtual derivative points $\mathbf{y}_{\mathrm{m}}$, hyper-hyperparameters $\zeta$
2:Solve (46) or (47) for $\bar{\Theta}$, where $\delta_{\mathrm{j}}:=\tilde{\boldsymbol{\mu}}_{\mathrm{j}}-\mathbf{m}_{\mathrm{j}}$, $\Lambda_{\mathrm{j}}(\Theta):=K_{\mathrm{j}}+\tilde{\Sigma}_{\mathrm{j}}$ defined in (30) and (57), and $S_{\boldsymbol{\gamma}}$ is given by (45).
3:Construct $\Gamma_{1}$, $\Gamma_{2}$ from $\gamma$.
4:Input: Prediction condition $\mathbf{u}$, $\mathbf{y}_{\mathrm{p}}$
5:Solve (41), (42), or (43) for $\mathbf{y}_{\mathrm{f}}$, where $m_{\mathrm{pm}}(\eta)$, $k_{\mathrm{pm}}(\eta)$ are given by (33) and (34), respectively.
6:Output: Output prediction $\mathbf{y}_{\mathrm{f}}$
3.6 Recovering Nonlinearities
In addition to output prediction, it could also be of interest to recover the input and output nonlinearities $\psi(\cdot)$ and $\phi(\cdot)$. The output nonlinearity estimate will also be used in Section 4 for DDPC design.
To obtain $\Phi(\mathbf{y})$ at any vector of query points $\mathbf{y}$, consider the joint distribution:
@@ 3.6 Recovering Nonlinearities
$\kappa_{y\mathrm{j}}(\mathbf{y})=\mathrm{col}\left(\frac{\partial\mathbf{k}_{y}(y_{\mathrm{m},1},\mathbf{y})}{\partial y_{\mathrm{m},1}},\dots,\frac{\partial\mathbf{k}_{y}(y_{\mathrm{m},n_{m}},\mathbf{y})}{\partial y_{\mathrm{m},n_{m}}},\kappa_{y}(\boldsymbol{\eta}^{d},\mathbf{y})\right).$
Similar to the derivation of (62) and (63), this leads to
$\Phi(\mathbf{y})|\mathcal{M},\boldsymbol{\chi}^{d}\sim\mathcal{N}\left(\mathbf{m}_{y\mathrm{p}}(\mathbf{y}),\mathbf{k}_{y\mathrm{p}}(\mathbf{y})\right),$ (48)
where
$\mathbf{m}_{y\mathrm{p}}(\mathbf{y})$ $:=\mathbf{m}_{y}(\mathbf{y})+\kappa^{\top}_{y\mathrm{j}}(\mathbf{y})\left(K_{\mathrm{j}}+\tilde{\Sigma}_{\mathrm{j}}\right)^{-1}\left(\tilde{\boldsymbol{\mu}}_{\mathrm{j}}-\mathbf{m}_{\mathrm{j}}\right),$ (49)
$\mathbf{k}_{y\mathrm{p}}(\mathbf{y})$ $:=\mathbf{k}_{y}(\mathbf{y},\mathbf{y})-\kappa^{\top}_{y\mathrm{j}}(\mathbf{y})\left(K_{\mathrm{j}}+\tilde{\Sigma}_{\mathrm{j}}\right)^{-1}\kappa_{y\mathrm{j}}(\mathbf{y}).$ (50)
Similarly, for $\Psi(\mathbf{u})$, we have
$\Psi(\mathbf{u})|\boldsymbol{\chi}^{d}\sim\mathcal{N}\left(\mathbf{m}_{u\mathrm{p}}(\mathbf{u}),\mathbf{k}_{u\mathrm{p}}(\mathbf{u})\right)$ (51)
with $\mathbf{m}_{u\mathrm{p}}(\mathbf{u})$ and $\mathbf{k}_{u\mathrm{p}}(\mathbf{u})$ similarly defined and $\kappa_{u\mathrm{j}}(\mathbf{u})=\mathrm{col}\left(\mathbf{0},\kappa_{u}(\boldsymbol{\eta}^{d},\mathbf{u})\right)$. Then, $\Psi(\mathbf{u})$ and $\Phi(\mathbf{y})$ are estimated by their posterior means $\mathbf{m}_{u\mathrm{p}}(\mathbf{u})$ and $\mathbf{m}_{y\mathrm{p}}(\mathbf{y})$, respectively.
4 Data-Driven Predictive Control with Implicit Gaussian Process Model
This section applies receding horizon DDPC using the implicit GP predictor described in Section 3. In detail, considering the problem of tracking an output reference $r_{k}$ at time $k$, the following expected control cost is considered at time $k$:
$J_{\mathrm{ctr}}:=\left\lVert\mathbf{u}^{k}\right\rVert_{R}^{2}+\mathbb{E}\left[\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{r}^{k})\right\rVert_{Q}^{2}\right],$ (52)
where $R\in\mathbb{R}^{n_{u}L^{\prime}\times n_{u}L^{\prime}}$, $Q\in\mathbb{R}^{n_{y}L^{\prime}\times n_{y}L^{\prime}}$ are the input and the output cost matrices, respectively, $\mathbf{r}^{k}:=\mathrm{col}\left(r_{k},\dots,r_{k+L^{\prime}-1}\right)$, $\mathbf{u}^{k}:=\mathrm{col}\left(u_{k},\dots,u_{k+L^{\prime}-1}\right)$, $\mathbf{y}^{k}_{0}:=\mathrm{col}\left(y_{k,0},\dots,y_{k+L^{\prime}-1,0}\right)$ are the reference, input, and true noise-free output trajectories at time $k$, respectively, and $L^{\prime}$ is the prediction horizon. In addition, we also consider input constraints $\mathbf{u}^{k}\in\mathcal{U}_{k}$ and linear output constraints $H\mathbf{y}^{k}_{0}\leq q$, where $H\in\mathbb{R}^{n_{c}\times n_{y}L^{\prime}}$ and $q\in\mathbb{R}^{n_{c}}$. Since the prediction error is unbounded, the output constraint can only be enforced with high probability $p$ as a chance constraint $\mathrm{Pr}\left(H\mathbf{y}^{k}_{0}\leq q\right)\geq p$.
In this section, the following continuity assumption on the output nonlinearity is considered.
Assumption 2.
@@ 4 Data-Driven Predictive Control with Implicit Gaussian Process Model
As in standard predictive control, an explicit output prediction of $\mathbf{y}^{k}_{0}$ could be obtained by applying Algorithm 2 to a particular prediction condition $\mathbf{u}=\mathrm{col}\left(u_{k-L_{0}},\dots,u_{k-1},\mathbf{u}^{k}\right),\ \mathbf{y}_{\mathrm{p}}=\mathrm{col}\left(y_{k-L_{0}},\dots,y_{k-1}\right)$. However, this requires solving (41), (42), or (43) as an inner problem, which is computationally challenging. Instead, we directly utilize the implicit predictor by quantifying the prediction error for an arbitrary input-output trajectory
@@
From (38), the distribution of $\Phi(\mathbf{y}^{k}_{0})$ is given by
$\Phi(\mathbf{y}^{k}_{0})\sim\mathcal{N}\left(\Phi(\mathbf{y}^{k})+m_{\mathrm{pm}}(\eta^{k}),\hat{k}_{\mathrm{pm}}(\eta^{k})\right),$ (53)
so the expected control cost can be calculated as
$J_{\mathrm{ctr}}=\left\lVert\mathbf{u}^{k}\right\rVert_{R}^{2}+\left\lVert\Phi(\mathbf{y}^{k})+m_{\mathrm{pm}}(\eta^{k})-\Phi(\mathbf{r}^{k})\right\rVert_{Q}^{2}\\ +\mathrm{tr}\left(Q\hat{k}_{\mathrm{pm}}(\eta^{k})\right).$
Since the output nonlinearity $\Phi(\cdot)$ is not known exactly, $\Phi(\mathbf{y}^{k})$ and $\Phi(\mathbf{r}^{k})$ are replaced by their posterior means $\mathbf{m}_{y\mathrm{p}}(\mathbf{y}^{k})$ and $\mathbf{m}_{y\mathrm{p}}(\mathbf{r}^{k})$ with certainty equivalence.
Then, the control problem optimizes $\mathbf{u}^{k}$ and $\mathbf{y}^{k}$ at the same time without considering the dependence of $\mathbf{u}^{k}$ and $\mathbf{y}^{k}$ explicitly.
Remark 11.
In standard GP-MPC, a one-step-ahead GP predictor is applied recursively, whereas a multi-step-ahead predictor is directly applied here. Multi-step-ahead predictors generally induce less conservatism in MPC and lead to less complex problems [20]. In particular, for GP-MPC, our approach circumvents the difficult problem of uncertainty propagation since the posterior of a GP from an input distribution is generally intractable, and the resulting distribution is not Gaussian [15].
Finally, the chance constraint satisfaction is guaranteed by the following lemma.
Lemma 2.
Under Assumptions 1 and 2, the chance constraint $\mathrm{Pr}\left(H\mathbf{y}^{k}_{0}\leq q\right)\geq p$ is satisfied if